# GlucOS: Security, correctness, and simplicity for automated insulin delivery

### v2

@@ Abstract
GlucOS combines algorithmic security, driver security, and end-to-end verification to protect against malicious ML models, vulnerable pump drivers, and drastic changes in human physiology. We use formal methods to prove correctness of critical components and incorporate humans as part of our defensive strategy. Our evaluation includes both a real-world deployment with seven individuals and results from simulation to show that our techniques generalize. Our results show that GlucOS maintains safety and improves glucose control even under attack conditions. This work demonstrates the potential for secure, personalized, automated healthcare systems.
1 Introduction
Type 1 Diabetes (T1D) is a metabolic disorder where an individual’s pancreas stops producing insulin. To compensate, they inject synthetic insulin. Computer systems, called automated insulin delivery systems, exist that use subcutaneous sensors to continuously monitor glucose concentrations (glucose is the body’s primary energy source) and inject insulin automatically using an insulin pump device (Figure 1). However, insulin is a dangerous hormone where too much insulin can kill people in a matter of hours [22] and too little insulin can kill people in a matter of days [13]. For automated insulin delivery systems, the key challenge is to maintain a balance: provide enough insulin to prevent dangerously high glucose levels while avoiding excessive insulin that could lead to life-threateningly low glucose levels. In this context, security means ensuring integrity for insulin dosing against malicious attacks or inadvertant errors. Several commercial [17, 5, 3] and open-source [42, 54] automated insulin delivery systems exist today, but none of them consider security as a primary design constraint.
@@ 1 Introduction
(c) Our system for automated insulin delivery.
@@
Figure 1: Overview of automated insulin delivery devices, closed-loop architecture, and a screenshot of GlucOS.
The complexity of automated insulin delivery systems and the fact that they make real-time healthcare decisions, makes their security paramount. Current research on security for these systems focuses primarily on evaluating algorithms using simulations [71, 72]. We argue that this approach falls short of addressing the full spectrum of challenges inherent in building trustworthy systems that operate within real-world constraints. Real-world systems must cope with runtime platform limitations, changes in human physiology, and the necessity of human interaction with the system. Plus, algorithms and simulations model attacks abstractly through their glucose response, missing out on details like which components of the system are likely to be compromised or how the system architecture can prevent damage from these potentially vulnerable components. Our experience shows that when faced with these practical challenges, algorithms designed in simulations often prove insufficient or inapplicable. These real-world complexities require a fundamentally different approach to designing security for automated insulin delivery systems. To address these challenges, we need practical systems that move beyond simulation-based designs and tackle the challenges of real-world systems.
@@
Current open-source automated insulin delivery systems are complex. For example, the OpenAPS [54] core function for calculating insulin doses consists of 1192 lines of Javascript code, 63 input and configuration parameters, and 90 branch statements. All of this logic influences the ultimate dosing decision, making reasoning about its correctness, both manually and automatically, challenging. Automated insulin delivery systems that use the OpenAPS algorithm embed it into native apps using a WebView for Javascript interpretation, adding to the overall complexity and increasing the attack surface. In contrast, our implementation of this same function is 25 lines of native and formally verified Swift code. None of the open-source automated insulin delivery systems use formal methods, can withstand attacks on components, or have mechanisms or policies for security.
@@
In this paper, we take on the challenge of building the first trustworthy automated insulin delivery system, called GlucOS. Our design includes: (1) an architecture where we apply separation principles for isolated and simple components, (2) a novel security mechanism and policy to enable secure ML-driven insulin delivery, (3) the ability to withstand attacks on insulin pump driver software, and (4) the application of formal methods to prove correct our insulin delivery path, the most sensitive part of the system. Despite these efforts, we identify scenarios in automated insulin delivery that require users to take action for security. Thus, we incorporate humans as part of our defensive strategy, providing them with agency and trust to take the required, corrective actions. Our approach is distinguished by its holistic consideration of the entire problem of automated insulin delivery. We provide clear security boundaries, mechanisms and policies to mitigate attacks on vulnerable components, and formal methods to provide assurances that our implementation is correct.
@@
Automated insulin delivery systems are different from traditional applications that the security community focuses on, but they are important. Research on operating systems or web browsers can list scores of CVEs from current systems and show how their system improves security by eliminating the impact of some or all of these vulnerabilities by virtue of their design. Automated insulin delivery systems do not have a long list of CVEs that we can reference, although there are a few [46, 47, 48, 49]. However, our position is that we need to make sure that these systems never get CVEs, because if they do, security vulnerabilities can kill people. Thus, we focus on building a new system, using security first principles and formal methods, to avoid security vulnerabilities and mitigate their damage from the start.
@@
  • The application of formal methods to prove correct the most critical parts of GlucOS.
  • A biological invariant that combines end-to-end physiological security checks with formal methods.
  • A case study describing our experiences from a real-world deployment with seven individuals, and results from simulation to show that our techniques generalize.
  • A human-centric design that considers the user as an integral part of the system, influencing both the overall architecture and including them in our defensive strategy.
We give top priority to the health, safety, and ethics of the humans using GlucOS to manage their T1D. Section 9.1 summarizes our safety protocol and Appendix A provides a detailed description.
There are 8.4 million people living with T1D worldwide [30], underscoring the need for providing trustworthy systems for automated insulin delivery. Our work, carried out with rigorous medical and ethical oversight, is a first step in addressing this need.
2 Overview
This paper describes our design for GlucOS, a system for trustworthy automated insulin delivery; we have two primary goals. First, we want the software to be simple and correct. Second, we want to identify the most vulnerable parts of the system and be able to prevent attacks on these components or withstand successful attacks. This section describes our overall architecture for GlucOS that strives to achieve these goals, and our threat model.
2.1 Locking down the insulin delivery path
@@
2.2 Architecture
From a system design perspective, we strive to keep our implementation simple and enable our use of formal methods to prove correct the most critical parts. For our software system architecture, we use separation principles from the OS and microkernel areas [8, 33] applied to the application layer to push complexity away from the most critical parts of the system. This architecture is similar to secure web browsers [31, 67, 63, 59], which also apply OS separation principles at the application layer. By decomposing the overall automated insulin delivery system, we strive to keep our trusted computing base simple.
Figure 2: Overall architecture for GlucOS.
Figure 2 shows our overall system architecture. At the core of our overall architecture is the BioKernel. The BioKernel is the component that interacts with the CGM and insulin pump hardware, executes the closed-loop dosing algorithm, runs security and safety checks, and produces event logs that other components use to learn the state of the system.
@@ 2.2 Architecture
To withstand attacks, the BioKernel serves as a reference monitor [10] to enforce security policies on the predictive ML model’s pump commands and the pump driver’s low-level I/O. This reference monitor architecture ensures that we can withstand attacks from malicious predictive ML as well as malicious pump drivers without having to know the internals of how they operate.
@@
2.3 Threat model
The most serious health risk for people who use synthetic insulin is low glucose, or hypoglycemia. Hypoglycemia results from an overdose of insulin. Once an automated insulin delivery system has delivered insulin, there is no way to remove it from the body. If untreated, severe hypoglycemia can lead to loss of consciousness, seizure, or death in a matter of hours. Thus, our primary security objective is to reduce or eliminate hypoglycemia that results from an attack on the system. Quantitatively, our goal is to maintain the American Diabetes Association’s recommended time in hypoglycemia of less than 4% [12].
@@ 2.3 Threat model
We assume that the Swift type system is correct and that the underlying operating system, the CGM driver and hardware, and the pump hardware we use are free of attacks and their device communications maintain integrity. Thus, we do not consider attacks emerging from tampered CGM sensors or pump hardware in our threat model. These are important but complementary problems to our work, which focuses on the system architecture and algorithms for automated insulin delivery.
@@
We also consider malicious pump drivers. In our threat model, malicious pump drivers can issue arbitrary pump I/O operations to add, drop, or modify commands sent to the pump.
@@
3 Algorithmic security: Withstanding malicious ML
Currently, no automated insulin delivery systems use deep neural network ML. Some current systems use statistical methods, like linear regression or model predictive control [21, 66, 43, 58, 54]. The reason for using this more traditional form of ML is sound. It provides consistent, explainable results and has a physiological basis that people can use to vet its decisions. However, all of the recent advances in ML, like image classifiers and large language models, have come from deep neural networks, which researchers have shown also work well for predicting metabolic states for use in automated insulin delivery systems [73, 35, 70, 65, 74, 64]. Despite this promise, none of the current automated insulin delivery systems incorporate more advanced deep neural networks due to the black-box nature of deep neural networks and the potentially dire consequences of mispredictions – an automated insulin delivery system’s unchecked “hallucination” [1] could have lethal consequences if it delivers an inappropriate insulin dose [22, 13].
@@ 3.1 Mechanism fundamentals
Our security mechanism draws inspiration from previous work on combining simple and complex models for safety [61, 51]. However, our approach extends these ideas by providing a novel security mechanism specifically designed for insulin delivery models. Our key contributions lie in our insight into the security properties of insulin delivery algorithms and our principles for how to apply our mechanism to automated insulin delivery systems (Section 3.2).
Figure 3: Algorithmic security mechanism for GlucOS.
@@ 3.2 Model principles
Algorithmically, we use a PID controller [25] from feedback control for our reactive safe model, and we use a deep neural-network from the literature for meal predictions for our predictive ML model [52]. Although we use this model in our current implementation, our system supports any predictive model. PID controllers are used in automated insulin delivery systems [68, 38] and research predictive ML models for insulin delivery have been well covered in the literature [73, 35, 70, 65, 74, 64].
@@
4 Driver security: Withstanding malicious pump drivers
@@
Pump drivers provide abstractions for insulin delivery applications to communicate with insulin pumps. Much like traditional device drivers, they have complex implementations since they help accomplish a wide range of tasks including pump setup/teardown, managing pump UX, handling insulin delivery/suspension, managing pump errors, accurately tracking insulin doses, and so on. Different pump drivers use different implementations with varying amounts of customizations and features to carry out these tasks.
@@
While CVEs do not currently exist for pump drivers, there is strong possibility for bugs or vulnerabilities in their implementation given their complexity and the fact that they not been formally verified. Since drivers control insulin pumps, these vulnerabilities can harm people using them, even lead to the death. For example, consider a scenario where the user decides to inject 3 units (3U) of insulin using the BioKernel. The BioKernel would relay this as a command to the pump through abstractions provided by the pump driver. If the pump driver inadvertently or intentionally issues this command twice, 6U of insulin would be delivered to the user, which would put the user in danger by dosing too much insulin. We need a mechanism to verify that the pump driver is executing correctly.
Our security goals for GlucOS’s driver security are to protect users from inadvertent bugs in pump drivers and safeguard against potentially malicious pump drivers. To achieve these goals, we (1) introduce driver security logic in the BioKernel that interposes on pump driver-pump communications, (2) validate all commands against known high-level pump commands, and (3) detect and handle anomalies by stopping any added/modified commands and proactively handling dropping critical commands.
This approach allows people to choose between various pump drivers while maintaining security, balancing user flexibility and safety with automated insulin delivery. It also keeps our trusted computing base small, simplifying formal verification.
4.1 Driver security logic design
@@
Figure 4: We interpose on communications between the pump driver and the pump to handle untrusted pump drivers to deliver insulin.
Message Message
name description
Request insulin delivery
getStatus status from the pump
Insulin dose delivery
statusResponse information from the pump
Insulin dose delivery
errorResponse error from the pump
Set default rate of
setBasalSchedule insulin delivery
Set temporary rate
setTempBasal of insulin delivery
bolus Deliver instant insulin dose
cancelDelivery Cancel insulin delivery
Figure 5: Types of insulin related messages exchanged between between a pump driver and an insulin pump.
@@
Figure 4 provides an overview of our driver security logic. The BioKernel registers high-level commands, like “deliver 2U of insulin” with our driver security logic, and sends them to the pump driver. The pump driver converts these commands into I/O messages to program the pump. However, these I/O messages come back to the driver security logic, which checks to make sure that they are consistent with the previously registered high-level pump command, before forwarding them to the pump hardware.
We only validate commands for insulin delivery since they pose the biggest threats to the user. We do not validate communication pertaining to pump setup / UX since users can directly detect the respective manipulations. Figure 5 lists the different types of messages pertaining to insulin delivery exchanged between a pump driver and the OmniBLE insulin pump, which we use in our current implementation. We list other message types in Appendix B.
By interposing on the pump I/O message path, the driver security logic can ensure that the driver has not added, modified, or dropped any pump commands. Detecting added or modified commands is relatively straightforward. Our driver security logic in the BioKernel knows all high-level pump commands that it sends to the driver. Thus, we have a simple logic to confirm the corresponding I/O messages from the driver. The driver security logic stops any anomalous I/O messages before sending them to the pump.
Detecting dropped commands is more nuanced when the command that the driver drops is a command to cancel an ongoing insulin delivery. Users can cancel insulin delivery if they accidentally triggered an insulin delivery command or incorrectly set its parameters. Since excess insulin delivery is dangerous, the driver security logic automatically sends these cancel commands to the pump when it detects a dropped insulin cancellation command. There are two commands that deal with canceling insulin delivery: cancelDelivery and setting the suspend option on setTempBasal. The cancelDelivery command cancels a previous insulin delivery command (if the delivery is still ongoing) and the suspend option on the setTempBasal command suspends all insulin delivery until the user resumes delivery.
The most interesting lesson learned from our experience is in coping with dropped pump commands. In our original design, our goal was to ensure that our driver security logic only understood I/O commands and prevented any malicious I/O commands from going through. However, our driver security logic ended up needing to issue its own I/O commands to deal with dropped insulin delivery cancellation commands. This more proactive approach to driver security is different than previous approaches that also use specifications for runtime driver verification [69, 63, 44], which simply infer I/O semantics and do not issue their own I/O commands. In general, we avoid programming the pump directly from the driver security logic. But in this case, we decided to program the pump directly from the driver security logic for human safety.
When we detect a malicious driver, we transition the system into a safe manual mode where the individual needs to take over control of their own insulin delivery. They can still use the BioKernel to manage their insulin, but they would be unable to use the automated closed-loop algorithms. In our evaluation, our system had no false positives and detected all attacks (Section 8.2). See Appendix C for more details on how we maintain safety even when detecting an attack.
4.2 Driver security logic implementation
We focus our implementation on Loop’s OmniBLE [4] pump driver. OmniBLE allows iOS applications to communicate with Omnipod pumps over Bluetooth. We include OmniBLE’s cryptography and Bluetooth communication modules as part of the BioKernel and incorporate the reference monitor just before encrypting messages sent to the pump/ decrypting messages received from the pump.
OmniBLE does not incorporate automatic retry in its implementation. However, OmniBLE occasionally sends additional commands as part of its operation. For example, when sending a setTempBasal command, if OmniBLE’s internal states indicate that delivery is still in progress as part of a previous setTempBasal command, it first issues a cancelDelivery command to cancel that delivery. To remove ambiguity in terms of whether the cancelDelivery command was issued as part of setTempBasal, we enforce OmniBLE to always issue the cancelDelivery command.
5 End-to-end security: Biological invariant
A key novelty in our overall approach is our introduction of an end-to-end check that ensures the BioKernel’s glucose calculations against observations of the user’s metabolic state. This raises an interesting challenge for formal verification as it requires modeling the true biological state of the user – a parameter that is both unknown at runtime and distinct from measured parameters which are part of the software implementation. We refer to the true safety invariant we enforce on the user’s state as the biological invariant, and proving it is both a key part of our implementation and one of the key conclusions of our formal verification efforts.

### v3

@@ Abstract.
GlucOS combines algorithmic security, driver security, and end-to-end verification to protect against malicious ML models, vulnerable pump drivers, and drastic changes in human physiology. We use formal methods to prove correctness of critical components and incorporate humans as part of our defensive strategy. Our evaluation includes both a real-world deployment with seven individuals and results from simulation to show that our techniques generalize. Our results show that GlucOS maintains safety and improves glucose control even under attack conditions. This work demonstrates the potential for secure, personalized, automated healthcare systems. Our source code is open source.
1. Introduction
Type 1 Diabetes (T1D) is a metabolic disorder where an individual’s pancreas stops producing insulin. To compensate, they inject synthetic insulin. Computer systems, called automated insulin delivery systems, exist that use subcutaneous sensors to continuously monitor glucose concentrations (glucose is the body’s primary energy source) and inject insulin automatically using an insulin pump device (Figure 1). However, insulin is a dangerous hormone as too much insulin can kill people in a matter of hours (cryer2012severe, ) and too little insulin can kill people in a matter of days (benoit2018trends, ). For automated insulin delivery systems, the key challenge is to maintain a balance: provide enough insulin to prevent dangerously high glucose levels while avoiding excessive insulin that could lead to life-threateningly low glucose levels. In this context, security means ensuring integrity for insulin dosing against malicious attacks or inadvertant errors. Several commercial (beta_bionics, ; tandem_iq, ; omnipod, ) and open-source (loop, ; openaps, ) automated insulin delivery systems exist today, but none of them consider security as a primary design constraint.
@@ 1. Introduction
(c) Overall GlucOS architecture.
@@
Figure 1. Overview of automated insulin delivery devices, closed-loop architecture, and GlucOS.
@@
Current open-source automated insulin delivery systems are complex. For example, the OpenAPS (openaps, ) core function for calculating insulin doses consists of 1192 lines of Javascript code, 63 input and configuration parameters, and 90 branch statements. All of this logic influences the ultimate dosing decision, making reasoning about its correctness, both manually and automatically, challenging. Automated insulin delivery systems that use the OpenAPS algorithm embed it into native apps using a WebView for Javascript interpretation, adding to the overall complexity and increasing the attack surface. In contrast, our implementation of this same function is 25 lines of native and formally verified Swift code. None of the open-source automated insulin delivery systems uses formal methods, can withstand attacks on components, or has mechanisms or policies for security.
@@
In this paper, we take on the challenge of building the first trustworthy automated insulin delivery system, called GlucOS. Our design includes: (1) an architecture where we apply separation principles for isolated and simple components, (2) a novel security mechanism and policy to enable secure ML-driven insulin delivery, (3) the ability to withstand attacks on insulin pump driver software, and (4) the application of formal methods to prove correct our insulin delivery path, the most sensitive part of the system.
Despite these efforts, we identify scenarios in automated insulin delivery that require users to take action for security. Thus, we incorporate humans as part of our defensive strategy, providing them with agency and trust to take the required, corrective actions.
Our approach is distinguished by its holistic consideration of the entire problem of automated insulin delivery. We provide clear security boundaries, mechanisms and policies to mitigate attacks on vulnerable components, and formal methods to provide assurances that our implementation is correct.
@@
Automated insulin delivery systems are different from traditional applications that the systems and security communities focus on, but they are important. Research on operating systems or web browsers can list scores of CVEs from current systems and show how their system improves security by eliminating the impact of some or all of these vulnerabilities by virtue of their design. Automated insulin delivery systems do not have a long list of CVEs that we can reference, although there are a few (cve1, ; cve2, ; cve3, ; cve4, ). However, our position is that we need to make sure that these systems never get CVEs, because if they do, security vulnerabilities can kill people. Thus, we focus on building a new system, using security first principles and formal methods, to avoid security vulnerabilities and mitigate their damage from the start.
@@
  • A case study describing our experiences from a real-world deployment with seven people, and results from simulation to show that our techniques generalize.
  • A human-centric design that considers the user as an integral part of the system, influencing both the architecture and including them in our defensive strategy.
We give top priority to the health, safety, and ethics of the humans using GlucOS to manage their T1D. Section 10.1 summarizes our ethical considerations.
There are 8.4 million people living with T1D worldwide (t1d_numbers, ), underscoring the need for providing trustworthy systems for automated insulin delivery. Our cross-disciplinary work, carried out with rigorous medical and ethical oversight, is a first step in addressing this need.
2. Overview
This paper describes our design for GlucOS, a system for trustworthy automated insulin delivery; we have two primary goals. First, we want the software to be simple and correct. Second, we want to identify the most vulnerable parts of the system and design a system to prevent attacks on these components or withstand successful attacks. This section describes our overall architecture for GlucOS that strives to achieve these goals, and our threat model.
2.1. Locking down the insulin delivery path
@@
2.2. Architecture
From a system design perspective, we strive to keep our implementation simple and enable our use of formal methods to prove correct the most critical parts. For our software system architecture, we use separation principles from the OS and microkernel areas (accetta1986mach, ; hartig1997performance, ) applied to the application layer to push complexity away from the most critical parts of the system. This architecture is similar to secure web browsers (grier2008secure, ; wang2009multi, ; tang2010trust, ; reis2019site, ), which also apply OS separation principles at the application layer. Figure 1(c) shows how we decompose our automated insulin delivery system, where we strive to keep our trusted computing base simple. GlucOS includes a BioKernel that forms our core trusted computing base. The BioKernel logs events that other components can ingest to learn the state of the system. These other components include apps for training ML models, visualizing statistics, an AI agent for suggesting dosing during meals, and an app for replaying closed-loop algorithm execution. Together, these components make up the overall GlucOS system.
Figure 2. Overall architecture for the BioKernel.
Figure 2 shows our architecture for the BioKernel. The BioKernel is the component that interacts with the continuous glucose monitor (CGM) and insulin pump hardware, executes the closed-loop dosing algorithm, runs security and safety checks, and produces event logs that other components use to learn the state of the system.
@@ 2.2. Architecture
To withstand attacks, the BioKernel serves as a reference monitor (anderson1972computer, ) to enforce security policies on the predictive ML model’s pump commands and the pump driver’s low-level I/O. This reference monitor architecture ensures that we can withstand attacks from malicious predictive ML as well as malicious pump drivers without having to know the internals of how they operate.
@@
2.3. Threat model
The most serious health risk for people who use synthetic insulin is low glucose, or hypoglycemia. Hypoglycemia results from an overdose of insulin. Once an automated insulin delivery system has delivered insulin, there is no way to remove it from the body. If untreated, severe hypoglycemia can lead to loss of consciousness, seizure, or death in a matter of hours. Thus, our primary security objective is to reduce or eliminate hypoglycemia that results from an attack on the system. Quantitatively, our goal is to maintain the American Diabetes Association’s recommended time in hypoglycemia of less than 4% (AmericanDiabetesAssociation2021, ).
@@ 2.3. Threat model
We assume that the Swift type system is correct and that the underlying operating system, the CGM driver and hardware, and the pump hardware we use are free of attacks and their device communications maintain integrity. Thus, we do not consider attacks from tampered CGM sensors or pump hardware in our threat model. These are important but complementary problems to our work, which focuses on the system architecture and algorithms for automated insulin delivery.
@@
We also consider malicious pump drivers. In our threat model, malicious pump drivers can issue arbitrary pump I/O to add, drop, or modify commands sent to the pump.
@@
3. Algorithmic security: Withstanding malicious ML
Currently, no automated insulin delivery systems use deep neural network ML. Some current systems use statistical methods, like linear regression or model predictive control (physio2011, ; physio2019, ; lunze2013control, ; control2019, ; openaps, ). The reason for using this more traditional form of ML is sound. It provides consistent, explainable results and has a physiological basis that people can use to vet its decisions. However, all of the recent advances in ML, like image classifiers and large language models, have come from deep neural networks, which researchers have shown also work well for predicting metabolic states for use in automated insulin delivery systems (Zhu2022, ; Jacobs, ; Zhang2021, ; Josep2020, ; RL, ; Tyler2022, ). Despite this promise, none of the current automated insulin delivery systems incorporate more advanced deep neural networks due to the black-box nature of deep neural networks and the potentially dire consequences of mispredictions – an automated insulin delivery system’s unchecked “hallucination” (chatgpt_hallucination, ) could have lethal consequences if it delivers an inappropriate insulin dose (cryer2012severe, ; benoit2018trends, ).
@@ 3.1. Mechanism fundamentals
Our security mechanism draws inspiration from previous work on combining simple and complex models for safety (sha2001using, ; mohan2013s3a, ). However, our approach extends these ideas by providing a novel security mechanism specifically designed for insulin delivery models. Our key contributions lie in our insight into the security properties of insulin delivery algorithms and our principles for how to apply our mechanism to automated insulin delivery systems (Section 3.2).
Figure 3. Algorithmic security mechanism for GlucOS.
@@ 3.2. Model principles
Algorithmically, we use a PID controller (franklin2002feedback, ) from feedback control for our reactive safe model, and we use a deep neural-network from the literature for meal predictions for our predictive ML model (mosquera2023enabling, ). Although we use this model in our current implementation, our system supports any predictive model. PID controllers are used in automated insulin delivery systems (weinzimer2008fully, ; kushner2018data, ) and research predictive ML models for insulin delivery have been well covered in the literature (Zhu2022, ; Jacobs, ; Zhang2021, ; Josep2020, ; RL, ; Tyler2022, ).
@@
4. Driver security: Withstanding malicious pump drivers
@@
Pump drivers provide abstractions for insulin delivery applications to communicate with insulin pumps. Much like traditional device drivers, they have complex implementations since they help accomplish a wide range of tasks including pump setup/teardown, managing pump UX, handling insulin delivery/suspension, managing pump errors, accurately tracking insulin doses, and so on.
@@
Given this complexity, there is a strong possibility for bugs or vulnerabilities to be present in existing pump drivers. These vulnerabilities can harm or even lead to the death of individuals since pump drivers control insulin pumps. For example, consider a scenario where the user decides to inject 3 units (3U) of insulin using the BioKernel. The BioKernel would relay this to the pump through abstractions provided by the pump driver. If the pump driver inadvertently or intentionally issues this command twice, 6U would be delivered, which could endanger the user by dosing too much insulin.
Our security goals for GlucOS’s driver security are to protect users from inadvertent bugs in pump drivers and safeguard against potentially malicious pump drivers. To achieve these goals, we introduce a shim between the pump driver and the pump, which validates all communications between them. We design our driver security mechanism to interpose on existing pump driver-pump communication instead of directly programming the pump. This design ensures that our trusted computing base remains small and simple by not including the complexity of pump driver implementations, thereby enabling formal verification. This design also provides flexibility with altering or introducing new functionality to pump drivers.
The most interesting aspect of our driver security mechanism was having to overrule our principle of not programming the pump in certain cases in the interest of protecting the individual. These cases involve commands pertaining to canceling ongoing insulin delivery. We could not directly adopt the design of other device driver security mechanisms (ref_validation, ; tang2010trust, ; expressos, ) for our driver security mechanism since these cases are specific to insulin pumps.
We focus our security efforts on pump drivers because they are complex (14.5k LoC), multi-threaded, event-driven, and use two-way communication between the pump and the driver. In contrast, CGM drivers are simple (1.9k LoC) and only read sensor values from the device. In our future work, we will formally verify our CGM driver using the same techniques we use for the insulin delivery path (Section 7).
4.1. Driver security mechanism design
@@
Figure 4. We interpose on communications between the pump driver and the pump to handle untrusted pump drivers.
@@
Figure 4 provides an overview of our driver security mechanism. The BioKernel registers high-level commands, like “deliver 2U of insulin” with our driver security mechanism, and sends them to the pump driver. The pump driver converts these commands into I/O messages to program the pump. The driver security mechanism then checks to make sure that these I/O messages are consistent with the previously registered high-level pump commands before forwarding them to the pump hardware. We only validate commands for insulin delivery since they pose the biggest threats to the user. On detecting added, modified or dropped commands from the pump driver, the driver security mechanism rejects further commands and transitions the system to a manual mode where the user takes control.
Dropped commands pertaining to canceling ongoing insulin delivery requires separate consideration since excess insulin is dangerous and cannot be withdrawn once it has been delivered. Users will cancel insulin delivery if they accidentally issued an insulin delivery command or incorrectly set the insulin dosage. In cases where a malicious driver drops a cancel command, the driver security mechanism directly issues I/O messages to the pump to cancel insulin delivery before transitioning the pump to a manual mode. We make the tradeoff of increasing the size of our trusted computing base to handle these cases in the interest of protecting the user. Fortunately, there are only two commands to cancel insulin delivery in our current implementation, so the increase to our trusted computing base is small.
5. End-to-end security: Biological invariant